suppose that 3 cards are drawn from a well-shuffled deck of 52 cards. what is the probability that all 3 are red?
Answer by nerdybill(6948)

suppose that 3 cards are drawn from a well-shuffled deck of 52 cards. what is the probability that all 3 are red?
total number of red cards: 26
total number of cards in deck: 52
.
probability of 1st card: 26/52 = 1/2
probability of 2nd card: 25/51
probability of 3rd card: 24/50 = 12/25
.
probability of getting a 3 red cards:
1/2 * 25/51 * 12/25
1/2 * 1/51 * 12
1/51 * 6
6/51
.118
11.8%